OpenStudy (anonymous):

The concession stand sells hot dogs and drinks during their high school foot ball game. John bought 6 hot dogs and 4 drinks and paid $6.70. Jessica bought 4 hot dogs and 3 drinks and paid $4.65. What is the price of a hot dog?

OpenStudy (anonymous):

Let the price of price of a hot dog and a drink be x and y dollars respectively therefore, converting '6 hot dogs and 4 drinks and paid $6.70' into equation we find 6x+ 4y=6.70 -----(1 similarly, converting ' 4 hot dogs and 3 drinks and paid $4.65' into equation we find 4x+ 3y=4.65 -----(2 Now in order to solve the simultaneous eq we need to eliminate at least on variable, so w multiply eq 1 by 2 and eq 2 by 3 and then subtracting 1 from 2 we have 12x+ 8y=13.40 -----(1 12x+ 9y=13.95 -----(2 - - - ----------------------- 0-y= -0.55 i.e. y= 0.55 substituting y= 0.55 in eq 1 we find 6x+ 4*0.55=6.70 -----(1 6x+ 2.2=6.70 6x= 6.70-2.2 6x= 4.5 x= 4.5/6 x=0.75 Thus, the price of a hot dog is $0.75